Grade Upto college levelTrigonometry

if the inequality sin2x +a cosx +a2> 1+cosx holds for any x belongs to R then the largest negative integral value of a is

A) -4 B) -3 C) -2 D) -1

ANAM V. RANGA REDDY
14 Years agoGrade Upto college level

1 Answer

Jitender Singh
12 Years ago
Ans: -2
Sol:
\sin ^{2}x + a\cos x + a^{2} > 1 + \cos x
1 - \sin ^{2}x + (1-a)\cos x - a^{2} < 0
\cos ^{2}x + (1-a)\cos x - a^{2} < 0
This is an simple quadratic equation in\cos x;
f(x) = \cos ^{2}x + (1-a)\cos x - a^{2}
\Delta \geq 0
\Delta = (1-a)^{2} - 4(-a^{2})
\Delta = 5a^{2} - 2a + 1
So, \Delta \geq 0
for all values of a.
Solution for\cos x
\cos x = ((a-1) \pm (5a^{2} -2a + 1)^{1/2})/2
-1\leq \cos x\leq 1
-2\leq (a-1)\pm (5a^{2} -2a + 1)^{1/2}\leq 2
-1-a\leq \pm (5a^{2} - 2a + 1)^{1/2}\leq 3-a
Square; we have
a^{2} + 2a + 1\leq 5a^{2} - 2a + 1\leq a^{2} - 6a + 9
We have two equation from here
a(a-1)\geq 0 …..............(1)
(a+2)(a-1)\leq 0…...........(2)
Solution of (1) & (2) is
a \epsilon [-2,0]
So, max. negative integer value for a is -2.
Cheer!
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty